OpenStudy (poohbear):

two angles are supplementary. one of the angles has a measure that is 52 degrees larger than the other angle. what is the measures of these two angles?

hartnn (hartnn):

let the smaller angle be x so larger angle is x+52 and as you must know,supplementary angles add up to 180 so x+x+52=180 can u find x from here??

OpenStudy (poohbear):

But, 128 and 52 doesn't have a difference of 52.

hartnn (hartnn):

x is not 128 x+x+52=180 so 2x+52=180 2x=128 x=64 and other angle is x+52=64+52=116

OpenStudy (poohbear):

Ohhhhh. I get it, thanks.